OpenStudy (sboodhoo):

A tennis ball has a radius of 6.7 centimeters. What is the volume of a tennis ball? Use 3.14 for pi. Round your answer to the nearest hundredth. 157.48 cm3 1259.19 cm3 4188.79 cm3 9634.08 cm3

OpenStudy (sboodhoo):

@mathmale

OpenStudy (sboodhoo):

@emmagrace231

OpenStudy (sboodhoo):

@kxttxn

OpenStudy (sboodhoo):

@sky060902

OpenStudy (whpalmer4):

Do you know the formula for the volume of a sphere, given its radius?

OpenStudy (anonymous):

http://www.calculatorsoup.com/calculators/geometry-plane/circle.php this is a circle calculater u can use

OpenStudy (kxttxn):

Hello! You can find the volume of a sphere using V=4/3πr^3 So just insert the radius for "r" and you should be able to find your answer!

OpenStudy (whpalmer4):

@sky060902 that will not help because it does not calculate volume. And this is not a difficult problem to do without such a calculator.

OpenStudy (anonymous):

oh okay thanks for letting me know :)

OpenStudy (whpalmer4):

As @kxttxn said, the formula for the volume of a sphere given its radius is \[V = \frac{4}{3}\pi r^3\]or \[V = \frac{4}{3}(3.14)r^3\]given that the problem specifies the value of \(\pi\) to use (\(3.14\)). Plug in the value you were given and round the answer to the nearest hundredth. What do you get?